OpenStudy (anonymous):

Solve the equation 1 over 2x + 2 = 5 for x.

OpenStudy (anonymous):

no its set up the way i put it.

OpenStudy (anonymous):

i just divide by -9 right..

OpenStudy (anonymous):

is it \[\frac{1}{2}x+2=5\]??

OpenStudy (anonymous):

or \[\frac{1}{x+2}=5\]??

OpenStudy (anonymous):